How many moles of H₂0 will be required to make 19.2 moles of O₂?


1
Expert's answer
2021-01-22T06:23:49-0500






As shown in the picture above,


according to the Stoichiometry ,

"\dfrac{n_{H_{2}O}}{n_{O_{2}}} = \dfrac{x}{19.2} = \dfrac{2}{1} 
\\
\\
x = 38 . 4 mol"